OpenStudy (anonymous):

ok say i did a exam and i gota 44 out of 77 what will my grade be?

OpenStudy (anonymous):

a 57

jimthompson5910 (jim_thompson5910):

Grade as a percentage = [ (# correct)/(# total) ]* 100 Grade as a percentage = (44/77) * 100 Grade as a percentage = ???

OpenStudy (anonymous):

57% Divide 44 by 77

OpenStudy (anonymous):

44/77=0.57=57%.